Teuvo Company entered into two contracts on the same date with Rundblad Corporation. Teuvo has provided the following analysis of price and cost for the contracts:

       Contract A

      Contract B

Contract price

$100,000

$ 50,000

Cost of related goods

62,000

54,000

Gross profit (loss)

$ 38,000

($4,000)

You inquired about why Teuvo would have agreed to make a sale at a loss. The company’s controller indicated that the company felt the profit on Contract A was sufficient to justify selling at a loss on Contract B, and it viewed the two contracts, together, as sufficiently profitable. “We normally get $85,000 for the goods in Contract A and $60,000 for the goods in Contract B, so we were happy with the transactions, but this is how the customer wanted to structure them.”

Required:

  1. Should the two contracts be combined for purposes of applying the 5-step revenue recognition model?

  2. What amount of revenue should Teuvo associate with each of the contracts?

  3. When should revenue be recognized on each of the contracts?
Answer #1

1. Yes , they must be combined.

reason: The contracts are with the same customer and hence multiple contracts with the same customer must be combined.

2. Stand alone price for contract A- 80,000

Stand alone price for contract B - 60,000

However, A has been sold at a price higher than usual , hence the additional $15000 must be allocated in the ratio of the standalone rates i.e. 85 :60.

Goods from Contract A = 85,000 + 15000 * 85 / 145 = 100,000 + 8793 = 93793

Goods from contract B = 50,000 + 15000 * 60/145 = 50000 + 6207 = 56207

3. Revenue is recognized when control of goods has transferred to the customer.